# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

### v1

@@ 4.3 Reduction to $\mathfrak{gl}_{n}$
$\begin{split}&{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{s},z\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})=\\ &\quad=\sum_{i=1}^{\ell}\sum_{j=k}^{n}\sum_{\rm part}\frac{{\mathbb{T}}_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{k}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{k-1})\,g({\bar{t}}^{k}_{\scriptscriptstyle\rm II},z)}\\ &\qquad\times\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.35)
@@
The case $-n\leq\ell<k<0$ and ${\bar{t}}^{0}=\varnothing$. This case is more subtle because it corresponds to the off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$, symmetric to ${\mathbb{B}}({\bar{t}})$ in the sence of the paper [21]. If according to the recurrence relations (3.10) the off-shell Bethe vector ${\mathbb{B}}({\bar{t}})$ is built from the entries $T_{i,j}(z)$, $1\leq i<j\leq n$ of $\mathfrak{gl}_{n}$-type, the symmetric off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$ is built from the $\mathfrak{gl}_{n}$-type matrix entries ${\mathbb{T}}_{i,j}(z)$ described in the lemma 2.1. Let us consider the recurrence relations (4.21) at $-n\leq\ell<k<0$ and empty set ${\bar{t}}^{0}=\varnothing$ in more details.
If ${\bar{t}}^{0}=\varnothing$ then according to the first line in (4.17) $|{\bar{t}}^{0}_{\scriptscriptstyle\rm III}|=0=\Theta(j)+\Theta(j-1)$ which is possible only for strictly negative values of the index $j$. So the recurrence relation (4.21) becomes
$\begin{split}\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\Big{|}_{{\bar{t}}^{0}=\varnothing}&={\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{-k-1},\{{\bar{t}}^{s},z_{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1})=\\ &=\mu^{k}_{\ell}(z;{\bar{t}})^{-1}\ \sum_{i=-n}^{\ell}\sum_{j=k}^{-1}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{1}^{n-1})\,,\end{split}$ (4.36)
@@
$\psi_{\ell}(z;{\bar{t}})=\frac{g({\bar{t}}^{-\ell},z_{-l})}{g(z_{-l},{\bar{t}}^{-\ell-1})},\qquad\phi_{k}(z;{\bar{t}})=\frac{g(z_{-k-1},{\bar{t}}^{-k-1})}{g({\bar{t}}^{-k},z_{-k-1})}$ (4.38)
@@
${\ell^{\prime}}=n+1+\ell,\quad 1\leq{\ell^{\prime}}\leq n-2,\qquad{k^{\prime}}=n+1+k,\quad 2\leq{k^{\prime}}\leq n.$ (4.40)
@@
$\begin{split}\hat{\lambda}_{k^{\prime}}(u)&=\frac{1}{\lambda_{n+1-{k^{\prime}}}(u+c({k^{\prime}}-1))}\prod_{s=1}^{{k^{\prime}}-1}\frac{\lambda_{n+1-s}(u+cs)}{\lambda_{n+1-s}(u+c(s-1))},\\ \hat{\alpha}_{{k^{\prime}}}(u)&=\frac{\hat{\lambda}_{k^{\prime}}(u)}{\hat{\lambda}_{{k^{\prime}}+1}(u)}=\frac{\lambda_{n-{k^{\prime}}}(u+c{k^{\prime}})}{\lambda_{n+1-{k^{\prime}}}(u+c{k^{\prime}})}=\frac{1}{\alpha_{n-{k^{\prime}}}(u+c{k^{\prime}})}\,.\end{split}$ (4.44)
Then according to the relation (2.18) which connect the values of the functions $\lambda_{k}(u)$ for negative and positive values of the index $k$ and definition (4.42) one gets
@@
$\begin{split}&\frac{1}{\phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm II})}\prod_{s=-j}^{-k-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})=\\ &\qquad=\frac{(-1)^{|{\bar{\tau}}^{{j^{\prime}}-1}_{\scriptscriptstyle\rm II}|-|{\bar{\tau}}^{{j^{\prime}}}|}}{h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{{k^{\prime}}}_{\scriptscriptstyle\rm II},z_{n})}\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{{\scriptscriptstyle\rm III}}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.47)
@@
$\begin{split}&{\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{{\ell^{\prime}}-1},\{{\bar{\tau}}^{s},z_{n}\}_{\ell^{\prime}}^{{k^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{k^{\prime}}^{n-1})=\\ &=\sum_{{i^{\prime}}=1}^{\ell^{\prime}}\sum_{{j^{\prime}}={k^{\prime}}}^{n}\sum_{\rm part}\frac{T_{i,j}(z){(-1)^{{\sf h}_{i}+{\sf h}_{j}}}\cdot{\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{{i^{\prime}}-1},\{{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}\}_{{i^{\prime}}}^{{\ell^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{{\ell^{\prime}}}^{{k^{\prime}}-1},\{{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}\}_{{k^{\prime}}}^{{j^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{{j^{\prime}}}^{n-1})}{\hat{\lambda}_{k^{\prime}}(z_{n})\,g(z_{n},{\bar{\tau}}^{{\ell^{\prime}}-1}_{\scriptscriptstyle\rm II})\,h({\bar{\tau}}^{\ell^{\prime}},z_{n})\,h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{k^{\prime}}_{\scriptscriptstyle\rm II},z_{n})}\\ &\qquad\times\prod_{s={i^{\prime}}}^{{\ell^{\prime}}-1}\Omega^{R}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm I},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm III}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.48)
Due to the definition (2.21) and the relations (2.18) and (2.9), the generators ${\sf h}_{i}$ for $i<0$ are equal to $-{\sf h}_{-i}$ and according to the formulas (2.28) the action of the generators ${\sf h}_{i}$ for $-n\leq i\leq-1$ to the off-shell Bethe vectors is given by
${{\sf h}_{i}}\cdot{\mathbb{B}}({\bar{t}})=(|{\bar{t}}^{-i}|-|{\bar{t}}^{-i-1}|)\ {\mathbb{B}}({\bar{t}})=(|{\bar{t}}^{\,n+1-{i^{\prime}}}|-|{\bar{t}}^{\,n-{i^{\prime}}}|)\ {\mathbb{B}}({\bar{t}})=(|{\bar{\tau}}^{{i^{\prime}}-1}|-|{\bar{\tau}}^{{i^{\prime}}}|)\ {\mathbb{B}}({\bar{\tau}})\,.$ (4.49)
This explains the appearance of the operator $(-1)^{{\sf h}_{i}+{\sf h}_{j}}$ in the r.h.s. of (4.48). Due to (4.49) this operator generate the sign factors $(-1)^{|{\bar{\tau}}^{{i^{\prime}}-1}|-|{\bar{\tau}}^{{i^{\prime}}}_{\scriptscriptstyle\rm II}|}$ and $(-1)^{|{\bar{\tau}}^{{j^{\prime}}-1}_{\scriptscriptstyle\rm II}|-|{\bar{\tau}}^{{j^{\prime}}}|}$ which comes from (4.46) and (4.47) respectively.
@@
Let us define the monodromy matrix entries ${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z)$ for all ${i^{\prime}},{j^{\prime}}=1,\ldots,n$ as:
@@
${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z_{n})=-T_{i,j}(z)\ (-1)^{\delta_{i,j}+{{\sf h}_{i}+{\sf h}_{j}}},\qquad\text{with }\begin{cases}i={i^{\prime}}-n-1\,,\\ j={j^{\prime}}-n-1\,.\end{cases}$ (4.50)
Lemma 2.1 ensures that this monodromy matrix is of $\mathfrak{gl}_{n}$-type and satisfies the $RTT$ relation (2.4) with the $\mathfrak{gl}_{n}$-type $R$-matrix. The off-shell Bethe vectors $\hat{\mathbb{B}}({\bar{t}})$ which can be constructed by the $\mathfrak{gl}_{n}$-type monodromy (4.50) and the recurrence relations (4.48) are related to the original vectors ${\mathbb{B}}({\bar{t}})$ as follows [21]
$\hat{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{n-1})=(-1)^{|{\bar{\tau}}|}\ {\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{n-1})=(-1)^{|{\bar{t}}|}\ {\mathbb{B}}(\varnothing,\{{\bar{t}}^{n-s}-c\,s\}_{1}^{n-1}),$ (4.51)
where $|{\bar{t}}|=|{\bar{\tau}}|$ is a total cardinality of the sets of Bethe parameters
$|{\bar{t}}|=\sum_{s=0}^{n-1}|{\bar{t}}^{s}|\,.$ (4.52)
The sign factor $(-1)^{|{\bar{t}}|}$ appears in the relation (4.51) because of the sign we introduced in definition (4.50) of the monodromy matrix ${\mathbb{T}}(z)$.
One can see now that the recurrence relation (4.36) or (4.48) for the Bethe vectors ${\mathbb{B}}({\bar{t}})$ coincides exactly with the recurrence relations (4.35) for the Bethe vectors $\hat{\mathbb{B}}({\bar{t}})$ if we remove the primes of the positive integers ${\ell^{\prime}},{k^{\prime}},{i^{\prime}},{j^{\prime}}$, replace $z_{n}\to z$, ${\bar{\tau}}^{s}\to{\bar{t}}^{s}$, $\hat{\lambda}_{k}(z_{n})\to\lambda_{k}(z)$, ${\mathbb{T}}_{i,j}(z_{n})\to{\mathbb{T}}_{i,j}(z)$. This proves that the recurrence relation (4.36) for the original Bethe vectors ${\mathbb{B}}({\bar{t}})$ is equivalent to the recurrence relation (4.35) for the Bethe vector $\hat{\mathbb{B}}({\bar{t}})$ (4.51) built from the monodromy matrix entries ${\mathbb{T}}_{i,j}(z)$ (4.50).
@@
$\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{B}_{\mathfrak{o}}(\{{\bar{t}}\}_{0}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\,.$ (4.53)
@@
The proof of this proposition is analogous to the proof of the proposition 3.2. We consider the recurrence relations (4.21) for three different ranges of the indices $\ell$ and $k$: $-a\leq\ell<k\leq a$, $a+1\leq\ell<k\leq n$ and $-n\leq\ell<k\leq-a-1$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathbb{B}_{\mathfrak{o}}(\varnothing)=\mathbb{B}_{\mathfrak{gl}}(\varnothing)=|0\rangle$, the recurrence relations will prove the statement of the proposition by induction on the cardinalities.

### v2

@@ 4.3 Reduction to $\mathfrak{gl}_{n}$
$\begin{split}&{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{s},z\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})=\\ &\quad=\sum_{i=1}^{\ell}\sum_{j=k}^{n}\sum_{\rm part}\frac{{\mathbb{T}}_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{k}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{k-1})\,g({\bar{t}}^{k}_{\scriptscriptstyle\rm II},z)}\\ &\qquad\qquad\qquad\times\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.34)
@@
The case $-n\leq\ell<k<0$ and ${\bar{t}}^{0}=\varnothing$. This case is more subtle because it is related to the off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$, connected to ${\mathbb{B}}({\bar{t}})$ in the sense of the paper [22]. If the off-shell Bethe vector ${\mathbb{B}}({\bar{t}})$ is built from the entries ${\mathbb{T}}_{i,j}(z)=T_{i,j}(z)$ for $1\leq i<j\leq n$ of $\mathfrak{gl}_{n}$-type according to the embedding (2.13), the off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$ is built from the $\mathfrak{gl}_{n}$-type matrix entries $\hat{{\mathbb{T}}}_{i,j}(z)$ described by the embedding (2.14).
Let us consider the recurrence relations (4.21) at $-n\leq\ell<k<0$ and empty set ${\bar{t}}^{0}=\varnothing$ in more details. If ${\bar{t}}^{0}=\varnothing$ then according to the first line in (4.17) $|{\bar{t}}^{0}_{\scriptscriptstyle\rm III}|=0=\Theta(j)+\Theta(j-1)$ which is possible only for strictly negative values of the index $j$. So the recurrence relation (4.21) becomes
$\begin{split}\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\Big{|}_{{\bar{t}}^{0}=\varnothing}&={\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{-k-1},\{{\bar{t}}^{s},z_{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1})=\\ &=-\ \mu^{k}_{\ell}(z;{\bar{t}})^{-1}\ \sum_{i=-n}^{\ell}\sum_{j=k}^{-1}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{1}^{n-1})\,,\end{split}$ (4.35)
@@
$\psi_{\ell}(z;{\bar{t}})=\frac{g({\bar{t}}^{-\ell},z_{-\ell})}{g(z_{-\ell},{\bar{t}}^{-\ell-1})},\qquad\phi_{k}(z;{\bar{t}})=\frac{g(z_{-k-1},{\bar{t}}^{-k-1})}{g({\bar{t}}^{-k},z_{-k-1})}$ (4.37)
@@
${\ell^{\prime}}=n+1+\ell,\quad 1\leq{\ell^{\prime}}\leq n-1,\qquad{k^{\prime}}=n+1+k,\quad 2\leq{k^{\prime}}\leq n.$ (4.39)
@@
$\begin{split}\hat{\lambda}_{k^{\prime}}(u)&=\frac{1}{\lambda_{n+1-{k^{\prime}}}(u+c({k^{\prime}}-1))}\prod_{s=1}^{{k^{\prime}}-1}\frac{\lambda_{n+1-s}(u+cs)}{\lambda_{n+1-s}(u+c(s-1))},\\ \hat{\alpha}_{{k^{\prime}}}(u)&=\frac{\hat{\lambda}_{k^{\prime}}(u)}{\hat{\lambda}_{{k^{\prime}}+1}(u)}=\frac{\lambda_{n-{k^{\prime}}}(u+c{k^{\prime}})}{\lambda_{n+1-{k^{\prime}}}(u+c{k^{\prime}})}={\alpha_{n-{k^{\prime}}}(u+c{k^{\prime}})}\,.\end{split}$ (4.43)
Then according to the relation (2.17) which connects the values of the functions $\lambda_{k}(u)$ for negative and positive values of the index $k$ and definition (4.41) one gets
@@
$\begin{split}\frac{1}{\phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm II})}\prod_{s=-j}^{-k-1}&\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})=\\ &\qquad=\frac{(-1)^{|{\bar{\tau}}^{{j^{\prime}}-1}_{\scriptscriptstyle\rm II}|-|{\bar{\tau}}^{{j^{\prime}}}|}}{h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{{k^{\prime}}}_{\scriptscriptstyle\rm II},z_{n})}\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{{\scriptscriptstyle\rm III}}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.46)
@@
$\begin{split}&{\mathbb{B}}^{\prime}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{-k-1},\{{\bar{t}}^{s},z_{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1})=\\ &=\sum_{{i^{\prime}}=1}^{\ell^{\prime}}\sum_{{j^{\prime}}={k^{\prime}}}^{n}\sum_{\rm part}\frac{T_{i,j}(z)\cdot{\mathbb{B}}^{\prime}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{-j-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{-j}^{-k-1},\{{\bar{t}}^{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{-\ell}^{-i-1},\{{\bar{t}}^{s}\}_{-i}^{n-1})}{\hat{\lambda}_{k^{\prime}}(z_{n})\,g(z_{n},{\bar{\tau}}^{{\ell^{\prime}}-1}_{\scriptscriptstyle\rm II})\,h({\bar{\tau}}^{\ell^{\prime}},z_{n})\,h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{k^{\prime}}_{\scriptscriptstyle\rm II},z_{n})}\\ &\qquad\qquad\qquad\times\prod_{s={i^{\prime}}}^{{\ell^{\prime}}-1}\Omega^{R}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm I},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm III}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.47)
where
${\mathbb{B}}^{\prime}({\bar{t}})=(-1)^{\sum_{s=1}^{n-2}|{\bar{t}}^{s}||{\bar{t}}^{s+1}|+\sum_{s=1}^{n-1}|{\bar{t}}^{s}|}{\mathbb{B}}({\bar{t}})\,.$ (4.48)
@@
Let us define the monodromy matrix entries $\hat{{\mathbb{T}}}_{{i^{\prime}},{j^{\prime}}}(z)$ for all ${i^{\prime}},{j^{\prime}}=1,\ldots,n$ defined by embedding (2.14) together with a shift of the spectral parameter as:
@@
$\hat{{\mathbb{T}}}_{{i^{\prime}},{j^{\prime}}}(z_{n})=T_{i,j}(z)\ ,\qquad\text{with}\quad\begin{cases}i={i^{\prime}}-n-1\,,\\ j={j^{\prime}}-n-1\,.\end{cases}$ (4.49)
The corresponding eigenvalues of $\hat{{\mathbb{T}}}_{k^{\prime}k^{\prime}}(z_{n})$ on the vacuum vector $|0\rangle$ coincide with $\hat{\lambda}_{k^{\prime}}(z_{n})$ from (4.44).
Note that the coefficients in the recursion relation (4.47) have exactly the same form as (4.34), if one replaces $z_{n}\to z$, ${\bar{\tau}}^{s}\to{\bar{t}}^{s}$, $\hat{\lambda}_{k^{\prime}}(z_{n})\to\lambda_{k^{\prime}}(z)$, $\hat{{\mathbb{T}}}_{i^{\prime},j^{\prime}}(z_{n})\to{\mathbb{T}}_{i^{\prime},j^{\prime}}(z)$.
Thus, the Bethe vector ${\mathbb{B}}({\bar{t}})$ relates to Bethe vectors $\hat{\mathbb{B}}({\bar{t}})$ constructed from the monodromy matrix $\hat{{\mathbb{T}}}(z)$ (4.49) by relation
$\hat{\mathbb{B}}(\varnothing,\{{\bar{t}}^{\,n-s}-cs-c\kappa_{n}\}_{1}^{n-1})=(-1)^{\sum_{s=1}^{n-2}|{\bar{t}}^{s}||{\bar{t}}^{s+1}|+\sum_{s=1}^{n-1}|{\bar{t}}^{s}|}\ {\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{n-1}).$ (4.50)
Such relation was studied in the paper [22], where it was shown that the vector Bethe $\hat{\mathbb{B}}({\bar{t}})$ is constructed from an inverted and transposed $\mathfrak{gl}_{n}$-type monodromy matrix, i.e. $\hat{{\mathbb{T}}}(z_{n})=({\mathbb{T}}(z)^{-1})^{\rm t}$.
@@
$\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\,.$ (4.51)
@@
The proof of this proposition is analogous to the proof of the proposition 3.2. We consider the recurrence relations (4.21) for three different ranges of the indices $\ell$ and $k$: $-a\leq\ell<k\leq a$, $a+1\leq\ell<k\leq n$ and $-n\leq\ell<k\leq-a-1$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathcal{B}_{\mathfrak{o}}(\varnothing)=\mathcal{B}_{\mathfrak{gl}}(\varnothing)=1$, the recurrence relations will prove the statement of the proposition by induction on the cardinalities.